# The Hidden Cost of Contextual Sycophancy: an AI Literacy Intervention in Human–AI Collaboration

### 18 May 2026 v1

@@ 4.3 Effects of the Intervention on Sycophantic Alignment
However, using a stricter measure of alignment, whether incorrect items were reproduced at the same rank position, a binomial generalized linear model revealed a significant interaction, b = -1.344, SE = 0.522, z = -2.57, p = .010. Following the intervention, the assistant was substantially less likely to mirror users’ incorrect rankings at the same positions compared to the control condition (OR = 0.26, 95% CI [0.09, 0.73]). A similar pattern emerged for rank-order alignment. Spearman correlations between user and assistant rankings of (at least 2) shared non-gold items (n = 87) showed a significant interaction, b = -1.053, SE = 0.314, z = -3.36, p = .001, indicating that the assistant’s ranking of incorrect items became significantly less aligned with the user’s ranking following the intervention. Although the treatment group exhibited higher baseline alignment, b = 0.748, SE = 0.182, z = 4.11, p < .001, the intervention led to a substantial reduction in alignment over time relative to the control condition.
@@ 5 Discussion and Future Work
The intervention did not reduce contextual error propagation (i.e., copying incorrect items). Increasing AI literacy and prompting may not be sufficient on their own to eliminate content-level dependence of the model on users’ initial framing. As suggested previously, prompting mitigation can help, but does not eliminate LLM sycophancy, especially in multi-turn or socially ambiguous contexts [9, 4]. However, our results showed that the intervention significantly decreased stronger forms of alignment, such as positional mimicry. The assistant continued to incorporate user-provided information but became less likely to mirror incorrect rankings directly. Some apparent sycophantic behavior may stem from general generative biases rather than alignment alone. LLMs may reproduce user inputs because these shift the token distribution, increasing the likelihood of reusing salient context elements without explicit agreement or “pleasing” intent. This aligns with recent theoretical work showing that repetition and bias can emerge from underlying attention dynamics [6].
@@ 5.0.1 Acknowledgements
This work was supported by IDEAL project, no. 2024-1-IT02-KA220-HED-000251425 and AILA project, no. 2025-1-ES01-KA220-HED-000352489.

### 21 May 2026 v2

@@ 4.3 Effects of the Intervention on Sycophantic Alignment
However, using a stricter measure of alignment, whether incorrect items were reproduced at the same rank position, a binomial generalized linear model revealed a significant interaction, b = -1.344, SE = 0.522, z = -2.57, p = .010. Following the intervention, the assistant was substantially less likely to mirror users’ incorrect rankings at the same positions compared to the control condition (OR = 0.26, 95% CI [0.09, 0.73]). A similar pattern emerged for rank-order alignment. Spearman correlations between user and assistant rankings of (at least 2) shared non-gold items (n = 87) showed a significant interaction, b = -1.053, SE = 0.314, z = -3.36, p = .001, indicating that the assistant’s ranking of incorrect items became significantly less aligned with the user’s ranking following the intervention. Although the experimental group exhibited higher baseline alignment, b = 0.748, SE = 0.182, z = 4.11, p < .001, the intervention led to a substantial reduction in alignment over time relative to the control condition.
@@ 5 Discussion and Future Work
The intervention did not reduce contextual error propagation. Increasing AI literacy and prompting may not be sufficient on their own to eliminate content-level dependence of the model on users’ initial framing. However, our results showed that the intervention significantly decreased stronger forms of alignment, such as positional mimicry. The assistant continued to incorporate user-provided information but became less likely to mirror incorrect rankings directly. Some apparent sycophantic behavior may stem from general generative biases rather than alignment alone. LLMs may reproduce user inputs because these shift the token distribution, increasing the likelihood of reusing salient context elements without explicit agreement or “pleasing” intent. This aligns with recent work showing repetition and bias can emerge from underlying attention dynamics [5].
@@ 5.0.1 Acknowledgements
This work was supported by IDEAL project, no. 2024-1-IT02-KA220-HED-000251425, AILA project, no. 2025-1-ES01-KA220-HED-000352489, Italian Ministry of University and Research under Grant No. 2023-NAZ-0206, PsyFuture – Dipartimento di Eccellenza 2023-2027, and by Volkswagen Foundation OpenUp Grant Ref. 9E530 Developing an Artificial Social Childhood (ASC).